### v1

# Elastic Metastability: Thermally Activated Snap-through Transitions in Nanostructures

Abstract
The morphological snap-through transition of a bistable elastic structure typically requires external loading, unless it is miniaturized to the nanoscale, where thermal fluctuations play a significant role. We propose that metastability can emerge in microscopic elastic systems when coupled to a thermal environment, giving rise to thermally activated transitions between metastable configurations. Using atomistic simulations combined with rare event methods, we demonstrate that snap-through events can occur as well-defined thermally activated transitions in a geometrically constrained graphene nanoribbon. Well-tempered metadynamics is employed to determine the transition energetics and pathways. Notably, the temperature dependence of the transition rate constant is accurately described by generalized transition state theory in terms of the Landau free energy. This study introduces a theoretical framework for understanding elastic metastability, extends reaction rate theory to nanomechanical systems, and suggests strategies for designing temperature-responsive nanodevices.
preprint: APS/123-QED
Thermally activated transitions underlie a great variety of processes in physics, chemistry, and biology, with examples ranging from magnetization reversal [1, 2] and solid state diffusion [3] to chemical reactions [4, 5] and protein folding [6]. A common characteristic of these systems is the presence of a free energy landscape with metastable states, between which transitions are driven by thermal fluctuations. In the context of elastic systems, although thermal fluctuations are not accounted for in a macroscopic elastic continuum, they may play an essential role in the microscopic scenario. Yet to date, thermally activated morphological transitions have neither been observed nor predicted in microscopic elastic systems. Therefore, we propose that, when miniaturized to the nanoscale, certain bistable elastic structures can undergo morphological transitions via thermal activation.
A bistable elastic structure possesses two stable equilibrium states. A morphological snap-through transition occurs when such a structure rapidly switches from one equilibrium state to the other. These transitions allow for the generation of fast movements through the release of stored elastic energy and are manifested in various biological and engineered systems, including the Venus flytrap [7], hummingbird beaks [8], mechanical metamaterials [9], and nanoelectromechanical switches [10]. In macroscopic systems, a snap-through transition is typically realized either by applying an external load that overcomes the energy barrier, or by imposing boundary actuation that destabilizes one of the equilibrium states [11]. However, due to the presence of thermal fluctuations in microscopic systems, such external interventions may not be required for a snap-through transition to occur. If a nanoscale bistable structure exhibits an energy barrier that can be surmounted by thermal fluctuations on physically relevant timescales, snap-through transitions would occur spontaneously with an associated rate constant, rendering the system metastable. While previous computational studies have investigated the snap-through of buckled graphene membranes [12, 13], the role of thermal activation in driving such transitions at the nanoscale remains unexplored, because the systems studied were subject to prohibitively large energy barriers precluding any spontaneous transition. In this work, we consider a buckled graphene nanoribbon (GNR) under geometric confinement, in which the energy barrier can be modulated to permit thermal activation.
The GNR can serve as a nanoscale counterpart of elastic strips, which have been increasingly adopted as a canonical system for analyzing the dynamics and mechanisms of snap-through transitions in bistable structures [14, 15, 16, 17, 18, 19, 20, 21]. Here, we consider a zigzag GNR with a length of 15.23 nm and N = 16 zigzag chains, whose edges are hydrogen-terminated to eliminate intrinsic mechanical instabilities [22]. When the GNR is placed under appropriate geometric constraints, in the same manner as a buckled elastic strip with clamped boundary conditions [19], it buckles into one of two stable configurations. Given the large length and time scales involved in the processes of interest, molecular dynamics (MD) simulations with empirical interatomic potentials constitute the only viable and effective computational approach that retains atomistic resolution. Empirical interatomic potentials have been widely used to describe the mechanical and thermodynamic properties of two-dimensional materials [23]. Among them, the second-generation reactive empirical bond-order (REBO-II) potential [24] is one of the most commonly used for solid carbon and hydrocarbon systems, and has been shown to reliably capture the elasticity [25] and thermomechanical behaviors [26] of graphene. Accordingly, we perform MD simulations for the H-terminated GNR system described by the REBO-II potential to investigate its thermally activated snap-through transitions. Furthermore, rare event methods, including enhanced sampling methods and reaction rate theory, are employed to unravel the free energy landscape and transition kinetics.
Figure 1: (a) Representative configurations of the GNR corresponding to the natural state (top), the asymmetric unstable equilibrium ($\mathcal{A}$ state, middle), and the inverted state (bottom). The rigidly constrained atoms at the two ends impose a fixed end-shortening $\Delta L$ and a symmetric tangent angle $\alpha$ at the boundaries. (b) The bifurcation diagram obtained from the Euler–beam model, with solid lines indicating the natural state (violet upper branch) and the inverted state (blue lower branch), and the orange dashed line indicating the $\mathcal{A}$ state. Diamond markers represent the equilibrium averages of $w_{0}$ from MD simulations.
The GNR can be transformed into a bistable structure when subject to geometric constraint at the boundaries. To achieve this, the outermost seven rows of atoms at each end are rigidly held, leaving a flexible segment of length $L$ = 13.75 nm. A fixed end-shortening $\Delta L=0.08L$ is applied between the two ends, which are then rotated in opposite directions by a small adjustable angle $\alpha$, causing the GNR to adopt one of two mechanical stable states: a natural state and an inverted state, as shown in Fig. 1(a). The equilibrium configurations of these states can be analyzed with Euler–Bernoulli beam theory [14, 19, 21] in the limit of small transverse displacement $w(x,t)$,
$\rho bh\frac{\partial^{2}w}{\partial t^{2}}+B\frac{\partial^{4}w}{\partial x^{4}}+F\frac{\partial^{2}w}{\partial x^{2}}=0,$ (1)
where $\rho$, $b$, and $h$ are the effective density, width, and thickness of the GNR, respectively; $B$ denotes the bending stiffness, and $F$ the applied compressive force. The geometric constraint enters the model through boundary conditions and the end-shortening confinement condition [14]. The time-independent solutions of Eq. (1) capture the equilibrium behavior of the system, which can be summarized in a bifurcation diagram in Fig. 1(b). The midpoint displacement $w_{0}$ of the GNR, weighted by $L$, is plotted as a function of the dimensionless control parameter $\mu=\alpha\sqrt{L/\Delta L}$. As $\mu$ increases, the inverted state loses its stability at $\mu^{*}=2$, where it intersects an asymmetric unstable equilibrium branch (denoted as $\mathcal{A}$ state) in a subcritical pitchfork bifurcation. In a macroscopic system, $\mu^{*}$ corresponds to the onset of snap-through towards the natural state. As the inverted state further approaches $\mu^{*}_{2}\approx 2.012$, it undergoes a saddle-node bifurcation and vanishes. Meanwhile, the natural state persists stably over the full range of $\mu$.
To compare the behavior of the GNR in atomistic simulations with the description of the Euler-beam equation, we monitor $w_{0}$ of the GNR at different values of $\mu$, by setting $\alpha$ to successive integer degrees. For each $\alpha$, the GNR is fully thermalized in the natural and the inverted states separately, and a 5 ns trajectory is initiated from each state to calculate the time average of $w_{0}$. A Langevin thermostat is used to provide the thermalizing environment at 300 K. The time-averaged $w_{0}$ from our MD simulations is found to be in good agreement with the Euler–beam model, as shown in Fig. 1(b). Naturally, the GNR configuration does not remain stationary at the mechanical equilibrium position of the natural or inverted state, but instead fluctuates around it, thereby defining a thermodynamic equilibrium state. When $\alpha$ exceeds $30^{\circ}$ ($\mu\approx 1.851$), the GNR cannot persist in the vicinity of the inverted state throughout a 5 ns trajectory—it snaps to the natural state under the influences of thermal fluctuations that are not captured by Eq. (1). This indicates that the thermal stability of the inverted state has already been compromised, even though the system is well below $\mu^{*}$, the threshold of mechanical instability.
Figure 2: (a) Midpoint displacement $w_{0}$ of the GNR (left y-axis, orange) and the corresponding atomistic potential energy (APE) (right y-axis, light blue) in a representative trajectory of the thermally activated snap-through transition. The zero of APE has been shifted for clarity. (b) Normalized time-correlation function of $w_{0}$ within the local equilibrium of the inverted state. (c) Cumulative distribution function of the escape times from the inverted state (blue solid line), compared with the maximum likelihood estimation fit (red dashed line).
Accordingly, we systematically investigate thermally activated snap-through transitions by tracking the snap-through events in 100 trajectories with different initial conditions sampled from an equilibrium ensemble of the inverted state at $\alpha=29^{\circ}$ ($\mu\approx 1.789$) and $T$ = 300 K. A representative transition trajectory is shown in Fig. 2(a). Prior to the snap-through event at $\approx$ 15 ns, the system stays in the local equilibrium of the inverted state, exhibiting small fluctuations of $w_{0}$. To characterize the relaxation of the system within the inverted state, we calculate the normalized time-correlation function of $w_{0}$, defined as
$\widetilde{C}(t)=\frac{\langle\delta w_{0}(t)\delta w_{0}(0)\rangle}{\langle\delta w_{0}(0)\delta w_{0}(0)\rangle},$ (2)
where $\delta w_{0}(t)=w_{0}(t)-\langle w_{0}\rangle$, and $\langle\cdot\rangle$ denotes the ensemble average in the inverted state. The profile of $\widetilde{C}(t)$ is presented in Fig. 2(b). The time integral of $\widetilde{C}(t)$ yields the dominant local relaxation time scale $\tau_{s}$ = 0.32 ns. In addition, we analyze the distribution of escape times in the 100 trajectories. The mean escape time is estimated using maximum likelihood estimation (MLE), yielding a value of $\tau_{e}=51.70$ ns, which is two orders of magnitude greater than the local relaxation time $\tau_{s}$. The cumulative distribution function (CDF) of the escape times is shown in Fig. 2(c), along with the fitted exponential distribution obtained from MLE. A Kolmogorov–Smirnov test produces a $p$-value of 0.957, confirming that the escape times follow an exponential distribution. The separation of time scales ($\tau_{e}\gg\tau_{s}$) [4, 5, 27] and the exponential distribution of escape times each demonstrate that the snap-through transitions in the GNR constitute a well-defined thermally activated rate process.
In Fig. 2(a), even while remaining within the local equilibrium states, the system exhibits substantial fluctuations in the atomistic potential energy, which obscures a clear description of the transition energetics. Therefore, instead of relying on elastic energy, a free-energy-based approach becomes necessary to capture the thermally activated snap-through at the nanoscale. Here, to determine the free energy profile of the system, we employ well-tempered metadynamics (WT-MetaD) [28], an enhanced sampling method that has been extensively used in complex molecular systems to accelerate rare events and accurately calculate free energy [29, 30].
Figure 3: (a) Free energy profile computed from WT-MetaD (five 500 ns production runs) along the order parameter $w_{0}$ for the GNR constrained at $\alpha=29^{\circ}$ ($\mu\approx 1.789$). Inset: Magnified view of the free energy profile highlighting the forward transition barrier $F^{+}$. (b) Forward and backward transition barriers, $F^{+}$ and $F^{-}$, as functions of the dimensionless geometric parameter $\mu$.
Because $w_{0}$ is the natural order parameter of the snap-through transition, one can perform one-dimensional WT-MetaD simulations using $w_{0}$ as the biased collective variable. At $\alpha=29^{\circ}$, the free energy profile $F(w_{0})$ computed from WT-MetaD is presented in Fig. 3(a), exhibiting two local minima corresponding to the inverted state and the natural state. The forward transition from the inverted to the natural state is associated with a free energy barrier of $F^{+}$ = 71.823 meV, while the backward transition corresponds to a much higher barrier of $F^{-}$ = 2526.776 meV. The substantial height of $F^{-}$ suggests that the reverse transition is kinetically suppressed, as will be addressed in our later analysis. Since the energetics of the system can be controlled by the geometric parameter $\mu$, we have further computed the free energy barriers from WT-MetaD at various values of $\mu$, as shown in Fig. 3(b). When $\mu=0$, the forward and backward transition barriers are equal, implying no energetic preference between the two states. As $\mu$ increases, the barriers become increasingly unequal.
Figure 4: Free energy surface computed from two-dimensional WT-MetaD in the ($w_{0}$, $\Delta w_{\text{qrt}}$) space. The collective variable $\Delta w_{\text{qrt}}$ quantifies the degree of asymmetry in the GNR. The locations of the natural state (red circle), inverted state (blue diamond), and the $\mathcal{A}$ state (yellow triangle), obtained from the Euler-beam model, coincide with the metastable and transition states identified on the FES. For illustrative clarity, the FES is shown for $\alpha=16^{\circ}$ ($\mu\approx 0.9873$).
To explore the free energy landscape of the system, we extend our approach to two-dimensional WT-MetaD involving two collective variables. Motivated by the important role of symmetry breaking in bistable elastic strips [19, 21], we propose a measure of asymmetry that is readily amenable to biased sampling in WT-MetaD, defined as the difference in transverse displacement between the quarter-length positions measured from the two ends of the GNR: $\Delta w_{\text{qrt}}=w(L/4)-w(-L/4)$. The WT-MetaD simulations are then performed by biasing both $w_{0}$ and $\Delta w_{\text{qrt}}$, resulting in the free energy surface (FES) in the ($w_{0}$, $\Delta w_{\text{qrt}}$) space, as shown in Fig. 4. Two basins corresponding to the inverted state and the natural state are identified, with local minima aligned with $\Delta w_{\text{qrt}}=0$, indicative of their symmetric nature. The transition pathways between the two states are characterized by the minimum free energy paths (MFEPs) connecting the basins, along which the system transitions between the two symmetric states via highly asymmetric intermediate states. To compare with the elastic continuum theory, the mechanical equilibrium positions from the Euler-beam model are plotted on the FES, showing good agreement with the metastable minima. Moreover, the asymmetric unstable equilibria ($\mathcal{A}$ state) coincide with a pair of first-order saddle points, indicating their role as the transition state in thermally activated transitions. It can be readily confirmed that the trajectories of one-dimensional WT-MetaD faithfully track the MFEPs on the two-dimensional landscape (see the Supplemental Material). Thus, the resulting free energy profile $F(w_{0})$ unambiguously corresponds to the Landau free energy governing the transitions.
Reaction rate theory has been widely employed to describe the kinetics of activated rate processes [4, 5], prominently including Kramers’ theory [31] and transition state theory (TST) [32]. Here we adopt generalized transition state theory in terms of the potential of mean force [33] to characterize the kinetics of the thermally activated snap-through transition, with emphasis on elucidating the temperature dependence of its transition rate constant.
Provided that there exists a collective variable $\xi$ such that the transition state dividing surface $\xi(\mathbf{r})=\xi^{\ddagger}$ separates the reactant region $\{\mathbf{r}|\xi(\mathbf{r})<\xi^{\ddagger}\}$ from the product region $\{\mathbf{r}|\xi(\mathbf{r})>\xi^{\ddagger}\}$ in configuration space, the transition rate constant can be calculated as the one-directional reactive flux through this surface. For a system characterized by the one-dimensional potential of mean force $W(\xi)$, the generalized TST rate constant takes the form
$k_{\mathrm{TST}}$ $=\frac{\langle{Z_{\xi}}^{\frac{1}{2}}\rangle_{\xi=\xi^{\ddagger}}}{(2\pi\beta)^{\frac{1}{2}}}\frac{\mathrm{e}^{-\beta W(\xi^{\ddagger})}}{\int_{\mathrm{R}}\mathrm{d}\xi\,\mathrm{e}^{-\beta W(\xi)}}$ (3)
$=\frac{\langle{Z_{\xi}}^{\frac{1}{2}}\rangle_{\xi=\xi^{\ddagger}}}{(2\pi\beta)^{\frac{1}{2}}{\Delta\xi}}\mathrm{e}^{-\beta\Delta F^{\ddagger}}.$ (4)
where $\langle\cdot\rangle_{\xi=\xi^{\ddagger}}$ denotes the ensemble average sample at the transition state, $\int_{\mathrm{R}}$ denotes the integral over the reactant region, and $\Delta\xi$ is given by $\Delta\xi=\int_{\mathrm{R}}\mathrm{d}\xi$; $\beta$ is the inverse temperature, $\Delta F^{\ddagger}$ is the free energy of activation, and $Z_{\xi}$ is the inverse effective mass associated with $\xi$, defined as
$Z_{\xi}\equiv\sum_{i=1}^{3N}\frac{1}{M_{i}}\left(\frac{\partial\xi}{\partial r_{i}}\right)^{2},$ (5)
with $M_{i}$ being the atomic mass associated with the $i$th component of $\mathbf{r}$. For the GNR system under investigation, the order parameter $w_{0}$ serves the role of $\xi$, so that the Landau free energy $F(w_{0})$ can be used in place of $W(\xi)$ in Eq. (3).
We justify the applicability of generalized TST to the activated snap-through transition by examining the consistency between the free energy of activation $\Delta F^{\ddagger}$ inferred from Eq. (4) based on the transition rates determined in unbiased simulations at different temperatures and the $\Delta F^{\ddagger}$ calculated directly from $F(w_{0})$ given by WT-MetaD. To this end, at $\alpha=29^{\circ}$, we first calculate $\Delta F^{\ddagger}$ in the temperature range from 300 K to 375 K by $\Delta F^{\ddagger}=F(w_{0}^{\ddagger})+k_{\mathrm{B}}T\ln{({\Delta w_{0}}^{-1}\int_{\mathrm{R}}\mathrm{d}w_{0}\,\mathrm{e}^{-\beta\Delta F(w_{0})})}$, where $\Delta w_{0}$ is the difference between the transition state value $w_{0}^{\ddagger}$ and the lower bound of $w_{0}$ sampled from the equilibrium ensemble of the inverted state. The resultant $\Delta F^{\ddagger}$ varies from 48.196 meV to 49.775 meV, showing that it is approximately constant over the temperature range considered.
Figure 5: (a) For the GNR system at $\alpha=29^{\circ}$, $\ln{(k\,T^{-\frac{1}{2}})}$ is plotted versus $1/T$ to reveal the temperature dependence of the transition rate constant $k$ computed from unbiased simulations. This plot is reminiscent of the Eyring plot [5], but features a different prefactor temperature dependence. Based on Eq. (4), $\Delta F^{\ddagger}$ can be inferred from the linear fit slope. (b) Mean escape time $\tau_{e}$ for the forward transition as a function of temperature, estimated using Eq. (3) for the system at $\alpha=24^{\circ}$.
In parallel, unbiased MD simulations are conducted to record the escape times for transitions from the inverted state to the natural state at 300 K, 325 K, 350 K, and 375 K. At each temperature, the transition rate constant $k$ is obtained as the inverse of the mean escape time $\tau_{e}$ from 100 independent trajectories (see the Supplemental Material). To analyze the temperature dependence of $k_{\mathrm{TST}}$, we recall that $w_{0}$ is computed as a linear combination of the midline C atom coordinates. This construction renders $\langle{Z_{w_{0}}}^{\frac{1}{2}}\rangle_{w_{0}={w_{0}}^{\ddagger}}$ effectively constant, leaving the temperature dependence of $k_{\mathrm{TST}}$ governed by $\beta$ in both the prefactor and the Boltzmann factor of Eq. (4). Motivated by this observation, we plot $\ln{(k\,T^{-\frac{1}{2}})}$ against $1/T$ in Fig. 5(a). The fit to the data exhibits good linearity, indicating that $k$ follows the temperature dependence described by Eq. (4). Moreover, the slope of the fit gives $\Delta F^{\ddagger}$ = 51.228 meV, in good agreement with the $\Delta F^{\ddagger}$ values previously calculated from WT-MetaD (average relative error $5\%$). These findings demonstrate that the temperature dependence of the rate constant $k$ is accurately captured by the generalized TST, and corroborate the self-consistency of our computational results. In practice, the actual rate constant $k$ generally deviates from the TST prediction $k_{\mathrm{TST}}$ by a multiplicative factor, the transmission coefficient $\kappa$, which accounts for recrossing effects. However, the consistent temperature dependence exhibited by $k$ and $k_{\mathrm{TST}}$ implies that, for our system, $\kappa$ remains approximately constant over at least a moderate range of temperature.
While generalized TST alone does not directly yield the absolute value of the rate constant $k$, Eqs. (3) and (4) still allow us to predict the kinetic behavior of the GNR system under various conditions by comparing relative changes in $\Delta F^{\ddagger}$ and temperature. In this context, we estimate the mean escape time $\tau_{e}$ for the forward transition at various temperatures, with the GNR constrained at $\alpha=24^{\circ}$ (Fig. 5(b)). A temperature variation of $\Delta T=20$ K within the studied range can result in orders-of-magnitude changes in $\tau_{e}$ and hence in the rate constant $k$. In addition, the backward transition is associated with $\Delta F_{\text{back}}^{\ddagger}=2.31$ eV at 300 K, corresponding to a mean escape time of $\sim 10^{23}$ years, which indicates that the natural state is effectively stable under ambient conditions. This analysis highlights that temperature control, along with adjusting the free energy via geometric constraints (Fig. 3(b)), can offer an effective strategy for tuning the transition rate, its temperature sensitivity, and the transition direction of the thermally activated snap-through in the GNR system.
In conclusion, we have shown that snap-through transitions in geometrically constrained GNRs can occur as well-defined thermally activated processes. By biasing the order parameter in WT-MetaD simulations, we have precisely characterized the free energy landscape, transition pathways, and the Landau free energy of the bistable GNR system. The temperature dependence of the transition rate constant is accurately captured by reaction rate theory. Our findings can be generalized to other bistable elastic nanostructures coupled to a thermalizing environment. With appropriately designed size, temperature, and structural energetics, such systems can exhibit morphological transitions via thermal activation at controllable transition rates. This study lays the foundation for understanding elastic metastability and opens up potential avenues for designing nanoscale thermal switches, thermal actuators, and other components in temperature-responsive devices.
This work has been supported by the National Natural Science Foundation of China under Grants No. 21673119 and No. 21973046.